### v2

# Separating the wave and particle attributes of two entangled photons

Abstract
We put forward a feasible scheme to spatially separate the wave and particle properties of two entangled photons by properly choosing the pre- and post-selection of path states. Our scheme, which implements the quantum Cheshire cat phenomenon for two-photon states, also guarantees that the observation of wave and particle properties of the two entangled photons always obey the Bohr’s complementarity principle.
pacs:
42.50.Wk,47.50.Ef, 42.50.−p, 03.67.−a
I Introduction
The wave-particle duality is one of the most characteristic phenomena featured by quantum mechanical systems. According to the Bohr’s complementarity principle (Bohr, 1928; Wootters and Zurek, 1979; Scully et al., 1991), one cannot simultaneously observe the wave and the particle behavior of a particle. In turn, quantum objects display its wave- or particle-like behavior depending on the experimental technique used to measure them (Wheeler, 1984; Grangier et al., 1986; Kurtsiefer et al., 1997; DÃŒrr et al., 1998; Marshall et al., 2003; Ionicioiu and Terno, 2011; Tang et al., 2012; Wang et al., 2019). The fundamental issues related to wave-particle duality and complementarity can be exemplified by Young-like double-slit experiments and their variants, e.g. Wheeler’s delayed choice experiment (Wheeler, 1978; Englert, 1996; Jacques et al., 2007, 2008; Ma et al., 2012; Manning et al., 2015). In those studies, the most important consequence of the wave-particle duality is the observation of quantum interference on a screen, independently on whether we send photons or particles onto the double-slit. Even more remarkable is the fact that quantum interference occurs also when one performs double-slit experiments with a single quantum object. A description of the experimental results using common language involves counterintuitive statements like "the particle passes through both slits at once", which, however, are not providing much insight. A more striking characterization of the phenomenon may be gained by means of the Wheeler’s delayed choice experiment.
In single-photon delayed-choice experiments (Jacques et al., 2007; Ionicioiu and Terno, 2011), a photon is sent into a Mach-Zehnder interferometer (MZI) where a balanced beam splitter (BS) splits it into two traveling paths. The term delayed choice is referring to the fact that one randomly determines whether to recombine the two paths or not, by a second BS, only after the photon has entered the MZI. If the second beam splitter is present, we observe interference fringes, indicating that the photon behaves as a wave, passing through both arms of the MZI. If BS2 is not there, each detector at the output fires with probability 1/2, i.e. only one of the two detectors is clicking at a time. This indicates that the photon traveled along a single arm i.e. it reveals a particle feature of the photon. With the double slit experiment, one realizes that photon has both a wave and a particle nature, and that these are inseparable. Experiments have shown that the above statements are true for any quantum systems including single electron, atom, neutron, and so on.
In general, since the wave and particle properties of a quantum object coexist simultaneously, one may assume that the two attributes cannot be separated from each other. However, recent works on the topic (Chowdhury et al., 2021; Li et al., 2023) has shown, upon exploiting the idea of quantum Cheshire cat (Aharonov et al., 2013), that one may spatially separates of wave and particle attributes of a quantum object. The Cheshire cat is a fictitious character in the famous novel “Alice in Wonderland”. The Cheshire cat has the ability of disappearing in space while its grin is still visible. In our usual life, we can imagine a cat without a grin, but it beyond our common sense to accept the existence of a grin without a cat as Alice said : “Well! I’ve often seen a cat without a grin, but a grin without a cat! It’s the most curious thing I ever saw in my life!”.
In 2013, Y. Aharonov and his collaborators introduced (Aharonov et al., 2013) a protocol to detach a quantum object from its inherent feature, and refer to the result as a quantum Cheshire cat. Quantum Cheshire cat may exist in a quantum system described by two-state vector formalism (Aharonov et al., 1964; Aharonov and Vaidman, 1990, 2008). Furthermore, its realization depends on performing a post-selected weak measurement (Aharonov et al., 1988) to extract the weak values (Aharonov and Botero, 2005) of the observables associated to the quantum object under investigation. In general, the weak value of an observable is a complex number and it can take arbitrary large values (Aharonov and Rohrlich, 2005). This feature of the weak value is usually referred to as an amplification effect of post-selection, and can be used to amplify tiny but useful information about physical systems (Hosten and Kwiat, 2008; Dixon et al., 2009; Starling et al., 2009, 2010; Pfeifer and Fischer, 2011; Egan and Stone, 2012; Zhou et al., 2013; Viza et al., 2013; Magaña Loaiza et al., 2014). Besides, weak measurement have been also employed in quantum state tomography (Lundeen et al., 2011; Lundeen and Bamber, 2012; de Gosson and de Gosson, 2012; Fischbach and Freyberger, 2012; Matzkin and Pan, 2013; Malik et al., 2013; Salvail et al., 2013; Di Lorenzo, 2013; Malik et al., 2014; Hariri et al., 2019; Pan et al., 2019), Hardy’s paradox (Aharonov et al., 2002; Jiang et al., 2018), three box paradox (Albert et al., 1985; Matzkin and Pan, 2013; Imran et al., 2021; Blasiak and Borsuk, 2022) and so on. For more details of weak measurement technique, we refer the reader to recent reviews of the field (Kofman et al., 2012; Svensson, 2013; Sokolovski, 2013; Dressel et al., 2014).
After the original QCC proposal (Aharonov et al., 2013), the phenomenon has been observed experimentally (Denkmayr et al., 2014) and received large attention theoretically (Guryanova et al., 2012; Matzkin and Pan, 2013; Lorenzo, 2013; Bancal, 2014; Ibnouhsein and Grinbaum, 2014; Corrêa et al., 2015; Duprey et al., 2018; Das and Pati, 2019, 2020; Das and Sen, 2021; Aharonov et al., 2021; Chowdhury et al., 2021; Hance et al., 2023a, b; Zhou et al., 2023; Ghoshal et al., 2023) and experimentally (Atherton et al., 2015; Ashby et al., 2016; Liu et al., 2020; Kim et al., 2021; Wagner et al., 2023; Danner et al., 2023; Li et al., 2023).
Among the different studies, in (Das and Pati, 2020; Liu et al., 2020) it has been proved that a quantum system may have a physical property erased and may obtain a new one that is did not initially have. More recently, a protocol to separate the wave and particle features of a single photon has been proposed (Chowdhury et al., 2021) and implemented (Li et al., 2023). Furthermore, in (Zhou et al., 2023), it has been investigated the possibility to spatially separate two entangled photons from their polarizations. Since entanglement is a crucial resource for quantum information processing and communication, a question arises on whether the separation of wave and particle attributes for entangled systems is possible. As we will see in the following, our results provide an affirmative answer to this question.
In this paper, motivated by previous studies (Chowdhury et al., 2021; Li et al., 2023; Zhou et al., 2023), we introduce a feasible theoretical model to spatially separate wave and particle properties of two entangled photons. After appropriate pre- and post-selection of the paths of photons, we can find the wave and particle features of the photons in different places. Besides their fundamental interest in the understanding of counterfactual communication, our results may promote the development of quantum metrology.
The remaining part of this paper is organized as follows. In Section II, we illustrate our motivations and briefly introduce the idea of quantum Cheshire cat. In SectionIII, we describe in details our proposal and show that by suitable pre- and post-selection one may realize the separation of wave and particle attributes of two entangled photons. In the same Section, we also describe a possible implementation of our scheme on an optical platform. Finally, in Section IV, we close the paper with some concluding remarks and outlooks.
II The quantum cheshire cat
In general, an object and its inherent properties always coexist and cannot be separated at anytime. Aharonov and collaborators found that the polarization (say, the grin) property can be separated from a photon (say, the cat) itself in a pre- and post-selected system. Here, we briefly summarize the ingredients required to generate such a quantum Cheshire cat. Initially, a horizontally polarized photon enters a Mach-Zehnder interferometer (MZI), and the pre-selected state of total system prepared is
$|\psi_{i}\rangle=\frac{1}{\sqrt{2}}\left(i|L\rangle+|R\rangle\right)|H\rangle,$ (1)
where $|L\rangle$ and $|R\rangle$ denote the left and right arms of MZI, i.e. the two possible paths of the photon, and $|H\rangle$ denotes the horizontal polarization of the photon. By adding and adjusting some optical elements, we may post-select the paths and polarization of the photon to obtain the following state vector
$|\psi_{f}\rangle=\frac{1}{\sqrt{2}}\left(|L\rangle|H\rangle+|R\rangle|V\rangle\right),$ (2)
where $|V\rangle$ represents the vertical polarization of the photon. Now, we measure the position (path) of the photon together with its circular polarization inside the MZI. In the two-state-vector formalism, the weak value of an observable $C$ is given by
$\langle C\rangle_{w}=\frac{\langle\psi_{f}|C|\psi_{i}\rangle}{\;\langle\psi_{f}|\psi_{i}\rangle}.$ (3)
The location of the photon, i.e, the findings of spatial degree of freedom of the photon, is described by the projectors
$\Pi_{\mu}=|\mu\rangle\langle\mu|,$ (4)
where $\mu\in\left(L,R\right)$, and $\Pi_{L}\left(\Pi_{R}\right)$ denote left (right) arm of MZI. On the other hand, the circular polarization of the photon in the left or right arms of MZI can be characterized by the operator,
$\sigma_{z}^{\mu}=\sigma_{z}\otimes\Pi_{\mu}.$ (5)
where $\sigma_{z}=|\uparrow_{y}\rangle\langle\uparrow_{y}|-|\downarrow_{y}\rangle\langle\downarrow_{y}|$ with $|\uparrow_{y}\rangle=\frac{1}{\sqrt{2}}\left(|H\rangle+i|V\rangle\right)$ and $|\downarrow_{y}\rangle=\frac{1}{\sqrt{2}}\left(|H\rangle-i|V\rangle\right)$, i.e. the eigenstate of the Pauli matrix $\sigma_{y}=\left(\begin{array}[]{cc}0&\ -i\\ i&0\end{array}\right)$ with corresponding eigenvalues $1$ and $-1$).
By taking into account the above pre- and post-selected states $|\psi_{i}\rangle$ and $|\psi_{f}\rangle$, the weak value of $\Pi_{\mu}$ and $\sigma_{z}^{\mu}$ are given by
$\langle\Pi_{\mu}\rangle_{w}=\delta_{\mu L},$ (6)
and
$\langle\sigma_{z}^{\mu}\rangle_{w}=\delta_{\mu R},$ (7)
respectively. Here, $\delta_{ij}$ denote the Kronecker $\delta$ function which defined as
$\delta_{ij}=\begin{cases}1,&i=j\\ 0.&i\neq j\end{cases}$ (8)
The weak values in Eq.(6) and Eq. (7) indicates that the spatial degree freedom of the photon is located in the left arm, while its circular polarization is detected with unit probability in the right arm of the MZI. This mean that in pre- and post-selected quantum systems one can realize the separation of the object from its inherent properties, as the Cheshire cat, whose smile may be discovered at a different place of the cat itself.
III Separation of wave-particle properties of two entangled photons
In this section, we describe our proposal for the separation of wave-particle properties of two-entangled photons. The schematic diagram of our proposed setup is shown in Fig. 1. It is based on the so-called WP toolbox introduced in (Rab et al., 2017) to convert the coherence superposition of the polarization states to the coherence superposition of the wave and particle entities.
Figure 1: Schematic diagram of the scheme to spatially separate the wave and particle properties of a two-photon entangled state.
In (Rab et al., 2017) single-photon states and two-photon entangled states have been analyzed. The generation processes of the wave-particle entangled state of two separated photons can be treated as parallel doubling of the single-photon scheme. In particular, upon using a half-wave plate (HWP, not shown in the figure) on the path of both the photons produced by spontaneous parametric down-conversion (SPDC), the two-photon state becomes
$|\psi\rangle_{in}=\cos\alpha|V\rangle|V\rangle+\sin\alpha|H\rangle|H\rangle,$ (9)
where $\alpha$ is an adjustable parameter, tunable by the HWP. After this stage, each photon enters one of two identical WP toolboxes, and the final output state may be written as (Rab et al., 2017)
$|\Phi\rangle_{out}=\cos\alpha|W\rangle|W^{\prime}\rangle+\sin\alpha|P\rangle|P^{\prime}\rangle.$ (10)
Here,
$|W\rangle$ $\equiv|Wave\rangle=e^{i\frac{\phi_{1}}{2}}\left(\cos\frac{\phi_{1}}{2}|1\rangle-i\sin\frac{\phi_{1}}{2}|3\rangle\right),$ (11)
$|W^{\prime}\rangle$ $\equiv\!\!|Wave^{\prime}\rangle=e^{i\frac{\phi_{1}}{2}}\left(\cos\frac{\phi_{1}}{2}|1^{\prime}\rangle\!-i\sin\frac{\phi_{1}}{2}|3^{\prime}\rangle\right),$ (12)
$|P\rangle$ $\equiv|Particle\rangle=\frac{1}{\sqrt{2}}\left(|2\rangle+e^{i\phi_{1}}|4\rangle\right),$ (13)
$|P^{\prime}\rangle$ $\equiv|Particle\rangle=\frac{1}{\sqrt{2}}\left(|2^{\prime}\rangle+e^{i\phi_{1}}|4^{\prime}\rangle\right).$ (14)
In the above expressions, $|n\rangle$ and $|n^{\prime}\rangle$ with $n,n^{\prime}\in\{1,2,3,4\}$ being the $n$-th output modes of from the two WP toolboxes (Rab et al., 2017), and $\phi_{1}$, $\phi_{1}^{\prime}$, $\phi_{2}$ and $\phi_{2}^{\prime}$ being controllable phase shifts in the two WP toolboxes. For more details about this generation of the states in Eqs. (11-14) we refer to the original paper (Rab et al., 2017). Notice that these states represent the capabilities ($|W\rangle$and $|W^{\prime}\rangle$) and incapabilities ($|P\rangle$ and $|P^{\prime}\rangle$) of the two photons to produce interference.
Finally, as shown in Fig. 1, one detects the two photons simultaneously by measuring the "wave operator" characterized by the projectors $X_{1}=|W\rangle\langle W|$ and $X_{2}=|W^{\prime}\rangle\langle W^{\prime}|$. The action of those operators is such that only the wave states $|W\rangle$ and $|W^{\prime}\rangle$ are transmitted whereas the particle states $|P\rangle$ and $|P^{\prime}\rangle$ are reflected. In this way, the state $|\Phi\rangle_{out}$ is transformed into
$|\psi_{i}\rangle$ $=\cos\alpha|R_{1}L_{2}\rangle|WW^{\prime}\rangle+\sin\alpha|L_{1}R_{2}\rangle|PP^{\prime}\rangle$
$=\cos\alpha|R_{1}L_{2}WW^{\prime}\rangle+\sin\alpha|L_{1}R_{2}PP^{\prime}\rangle.$ (15)
This is our pre-selected state. Here, as quantum Cheshire cat’s notation, we use the composite system of path and attributes of two photons as |..$|.\ .\rangle_{path}|.\ .\rangle_{attributes}$. $|L_{i}\rangle$ and $|R_{i}\rangle$ denoting the left and right paths , $i\in\{1,2\}$ labels the photons. $|R_{1}L_{2}\rangle|WW^{\prime}\rangle$ represents $1$st (2nd) photon’s wave property $|W\rangle$ ($|W^{\prime}\rangle$) at $|R_{1}\rangle$ ($|L_{2}\rangle$), and $|L_{1}R_{2}\rangle|PP{}^{\prime}\rangle$ represents $1$st (2nd) photon’s particle property $|P\rangle$ ($|P^{\prime}\rangle$) at $|L_{1}\rangle$ ($|R_{2}\rangle$). $|W\rangle$, $|W^{\prime}\rangle$ and $|P\rangle$, $|P^{\prime}\rangle$ denote, respectively, the wave and particle states as (Ionicioiu and Terno, 2011; Rab et al., 2017)
In order to realize the separation of wave and particle attributes of two entangled photons, the post-selected state should be of the form
$|\psi_{f}\rangle=\frac{1}{\sqrt{2}}\left(|R_{1}L_{2}WW^{\prime}\rangle+|L_{1}R_{2}PP^{\prime}\rangle\right).$ (16)
This means that in our scheme the post-selected state is a special case of the pre-selected class of states with $\alpha=\frac{\pi}{4}$.
The determination of the post-selected state $|\psi_{f}\rangle$ is crucial for the implementation of our scheme. As shown in Fig. 1, its verification can accomplished by an optical setup made of five beam splitters $BS_{i}$ ($i=1,2,3,4,5$), two $\sigma^{1224}$operators, the $X_{3}$ and $X_{4}$ operators, three mirrors $M_{i}$ ($i=1,2,3$), and six detectors $D_{i}$ ($i=1,2,3,4,5,6$). The post-selection occurs if and only if the detector $D_{5}$ clicks. On the contrary, if our post-seleceted state is $|\psi_{f}\rangle$, then, in the optical arrangement of Fig. 1, we can guarantee only $D_{5}$ give the answer ,“yes”. Let us prove this statement.
At first, the photons in $L_{1}$ and $R_{2}$ arms go through the $BS$ followed by $\sigma^{1234}$ operator. Here, effect of $\sigma^{1234}$ operator is to switch the photon paths in the four-dimensional basis $\{|1\rangle,|2\rangle,|3\rangle,|4\rangle\}$. Its matrix form can be written as (Rab et al., 2017)
[…]
and its base vectors can be written as
$|1\rangle$ $=(\begin{array}[]{cccc}1&0&0&0\end{array})^{T},|2\rangle=(\begin{array}[]{cccc}0&1&0&0\end{array})^{T},$ (18c)
$|3\rangle$ $=(\begin{array}[]{cccc}0&0&1&0\end{array})^{T},|4\rangle=(\begin{array}[]{cccc}0&0&0&1\end{array})^{T}.$ (18f)
The action of $\sigma^{1234}$ is to switch paths as follows (Li et al., 2023)
$|1\rangle\stackrel{{\scriptstyle[}}{{\sigma}}^{1234}]{\sigma^{1234}}{\rightleftharpoons}|2\rangle,\ |3\rangle\stackrel{{\scriptstyle[}}{{\sigma}}^{1234}]{\sigma^{1234}}{\rightleftharpoons}|4\rangle.$ (19)
The beam splitter $BS_{1}$ is chosen as to perform the transformations $|2\rangle\rightarrow\frac{1}{\sqrt{2}}(|2\rangle+|4\rangle)$, and $|4\rangle\rightarrow\frac{1}{\sqrt{2}}(|2\rangle-|4\rangle)$. In this way, the particle state can converted to the wave state after the action of $BS_{1}$ and $\sigma^{1234}$, i.e,
$|P\rangle\overset{BS_{1}}{\longrightarrow}$ $\frac{1}{\sqrt{2}}\left[\frac{1}{\sqrt{2}}(|2\rangle+|4\rangle)+e^{i\phi_{1}}\frac{1}{\sqrt{2}}(|2\rangle-|4\rangle)\right]$
$=$ $e^{i\frac{\phi_{1}}{2}}\left(\cos\frac{\phi_{1}}{2}|2\rangle-i\sin\frac{\phi_{1}}{2}|4\rangle\right)\overset{\sigma^{1234}}{\longrightarrow}$
$e^{i\frac{\phi_{1}}{2}}\left(\cos\frac{\phi_{1}}{2}|1\rangle-i\sin\frac{\phi_{1}}{2}|3\rangle\right)=|W\rangle.$ (20)
The same happens for primed paths with $|n^{\prime}\rangle\in\left(|1^{\prime}\rangle,|2^{\prime}\rangle,|3^{\prime}\rangle,|4^{\prime}\rangle\right)$. Thus, after the photons in $L_{1}$ and $R_{2}$ arms simultaneously go through beam splitters $BS_{1}$ and $BS_{2}$ followed by the $\sigma^{1234}$ operators, our post-selected state is given by
$|\psi_{f^{\prime}}\rangle=\frac{1}{\sqrt{2}}\left(|R_{1}L_{2}\rangle+|L_{1}R_{2}\rangle\right)|WW^{\prime}\rangle.$ (21)
Next, by adjusting the beam-splitters $BS_{3}$ and $BS_{4}$, respectively, we make the states $|R_{1}L_{2}\rangle$ and $|L_{1}R_{2}\rangle$ go through $X_{3}$ and $X_{4}$, such that the detectors $D_{1}$ and $D_{2}$ will not click. The $X_{3}=X_{4}=|W\rangle\langle W|$ operators transmit only the wave state, and this action guarantees that detectors $D_{3}$ and $D_{4}$ will not click. Finally, by adjusting the beam splitter $BS_{5}$, the state $|R_{1}L_{2}\rangle+|L_{1}R_{2}\rangle$ reaches the detector $D_{5}$. Overall, we conclude that if and only if the detector $D_{5}$ clicks we have made a successful post-selection. Otherwise, if any other detector clicks post-selection fails, and we fail to separate the wave and particle attributes of the two entangled photons.
The operators that measure whether the wave and particle attributes of two entangled photons are present in the left and right arms is defined as
$\Pi_{\nu}^{\mu_{i}}=|\mu_{i}\rangle\langle\mu_{i}|\otimes|\nu\rangle\langle\nu|,$ (22)
with $\nu\in\{W,W^{\prime},P,P^{\prime}\}$, $\mu_{i}\in\{L_{i},R_{i}\}$ and $i\in\{1,2\}$. After preparing the pre-selected state, we choose to post-select by $|\psi_{f}\rangle$ and perform weak measurement for observables $\Pi_{\nu}^{\mu_{i}}$. By substituting the above-defined pre- and post-selected states into Eq. (3), we find the corresponding weak values of the observables $\Pi_{\nu}^{\mu_{i}}$, i.e.,
$\langle\Pi_{W}^{\mu_{i}}\rangle$ $=\frac{\cos\alpha}{\cos\alpha+\sin\alpha}\delta_{1i}\delta_{R_{i}\mu_{i}},$ (23a)
$\langle\Pi_{P}^{\mu_{i}}\rangle$ $=\frac{\sin\alpha}{\cos\alpha+\sin\alpha}\delta_{1i}\delta_{L_{i}\mu_{i}},$ (23b)
$\langle\Pi_{W^{\prime}}^{\mu_{i}}\rangle$ $=\frac{\cos\alpha}{\cos\alpha+\sin\alpha}\delta_{2i}\delta_{L_{i}\mu_{i}},$ (23c)
$\langle\Pi_{P^{\prime}}^{\mu_{i}}\rangle$ $=\frac{\sin\alpha}{\cos\alpha+\sin\alpha}\delta_{2i}\delta_{R_{i}\mu_{i}}.$ (23d)
If $\alpha\neq 0,\frac{\pi}{2}$, these weak values indicated that the wave and particle attributes of two entangled photons are perfectly separated, and in each one of the four arms we only have one attribute (wave or particle ) of the two photons. In the particular case, $\alpha=\frac{\pi}{4},$ the weak values for the wave and particle attributes may be written as $\langle\Pi_{P}^{L_{1}}\rangle=\langle\Pi_{W}^{R_{1}}\rangle=\frac{1}{2}$, $\langle\Pi_{P^{\prime}}^{R_{2}}\rangle=\langle\Pi_{W^{\prime}}^{L_{2}}\rangle=\frac{1}{2}$. This means that for $\alpha=\frac{\pi}{4}$ case, half of the particle attributes of the two entangled photons occurrs in $L_{1}$ and $R_{2}$ arms and half of the wave attributes are present in $R_{1}$ and $L_{2}$ arms of the interferometer, respectively. Furthermore, it is clear that these weak values cannot take arbitrary large values, and always obey the Bohr’s complementarity principle, i.e.,
$\langle\Pi_{P}^{L_{1}}\rangle+\langle\Pi_{W}^{R_{1}}\rangle$ $=1,$ (24)
$\langle\Pi_{P^{\prime}}^{R_{2}}\rangle+\langle\Pi_{W^{\prime}}^{L_{2}}\rangle$ $=1.$ (25)
IV conclusion
In this paper, we have investigated the possibility of spatially separating the wave and particle features of two entangled photons. Our work is based on measuring the weak values of two-level quantum systems by selecting appropriate pre- and post-selection states in order to generate polarization-path like quantum Cheshire cats.
We found that properly pre- and post-selecting the two-photon state, the wave and particle attributes of two entangled photons can be spatially separated, while keeping the validity of Bohr’s complementarity principle. We also proposed a feasible setup for the implementation of our thought experiment on a quantum optical platform.
The quantum entanglement is a cornerstone of quantum informatics and it has an irreplaceable role in quantum precision measurements. In particular, in quantum metrology one would generally require that the measurement of a physical quantity on an entangled state does non disturb to other variable. Our results may foster the exploration of schemes to separate entangled observables and enhance precision of quantum metrological protocols. At the same time, in the emerging field of counterfactual communication (Salih et al., 2013; Li et al., 2015), information may be exchanged even without the transmission of quantum state. In this framework, entanglement also plays a vital role, but the basic principles of communication is still unclear. Also in this case, our proposal may promote the understanding of the fundamental mechanism behind counterfactual communication.
Finally, we anticipate that our theoretical proposal is feasible with current technology and may be implemented in quantum optical labs in the near future.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (No. 12365005).

### v4

@@
 Abstract
-We put forward a feasible scheme to spatially separate the wave and particle properties of two entangled photons by properly choosing the pre- and post-selection of path states.
-Our scheme, which implements the quantum Cheshire cat phenomenon for two-photon states, also guarantees that the observation of wave and particle properties of the two entangled photons always obey the Bohr’s complementarity principle.
+Wave-particle duality is one of the most intriguing counterfactual concepts in quantum theory.
+In our common sense, the wave and particle properties of a quantum object are inseparable.
+However, the recent studies based on Quantum Cheshire Cat phenomena showed that separating the physical properties of a quantum object including wave and particle attributes from itself are possible in microscopic system described by two-state vector formalism.
+In this study, we put forward a feasible scheme to spatially separate the wave and particle attributes of two entangled photons by properly choosing the pre- and post-selection of path states.
+Our scheme also guarantees that the observation of wave and particle properties of the two entangled photons always obey the Bohr’s complementarity principle.
 pacs:
-42.50.Wk,47.50.Ef, 42.50.−p, 03.67.−a
+03.65.-w, 03.65.Ta, 42.50.−p, 03.67.−a
 I Introduction
-The wave-particle duality is one of the most characteristic phenomena featured by quantum mechanical systems.
-According to the Bohr’s complementarity principle (Bohr, 1928; Wootters and Zurek, 1979; Scully et al., 1991), one cannot simultaneously observe the wave and the particle behavior of a particle.
-In turn, quantum objects display its wave- or particle-like behavior depending on the experimental technique used to measure them (Wheeler, 1984; Grangier et al., 1986; Kurtsiefer et al., 1997; DÃŒrr et al., 1998; Marshall et al., 2003; Ionicioiu and Terno, 2011; Tang et al., 2012; Wang et al., 2019).
-The fundamental issues related to wave-particle duality and complementarity can be exemplified by Young-like double-slit experiments and their variants, e.g.
-Wheeler’s delayed choice experiment (Wheeler, 1978; Englert, 1996; Jacques et al., 2007, 2008; Ma et al., 2012; Manning et al., 2015).
-In those studies, the most important consequence of the wave-particle duality is the observation of quantum interference on a screen, independently on whether we send photons or particles onto the double-slit.
-Even more remarkable is the fact that quantum interference occurs also when one performs double-slit experiments with a single quantum object.
-A description of the experimental results using common language involves counterintuitive statements like "the particle passes through both slits at once", which, however, are not providing much insight.
-A more striking characterization of the phenomenon may be gained by means of the Wheeler’s delayed choice experiment.
-In single-photon delayed-choice experiments (Jacques et al., 2007; Ionicioiu and Terno, 2011), a photon is sent into a Mach-Zehnder interferometer (MZI) where a balanced beam splitter (BS) splits it into two traveling paths.
-The term delayed choice is referring to the fact that one randomly determines whether to recombine the two paths or not, by a second BS, only after the photon has entered the MZI.
-If the second beam splitter is present, we observe interference fringes, indicating that the photon behaves as a wave, passing through both arms of the MZI.
-If BS2 is not there, each detector at the output fires with probability 1/2, i.e.
-only one of the two detectors is clicking at a time.
-This indicates that the photon traveled along a single arm i.e.
-it reveals a particle feature of the photon.
-With the double slit experiment, one realizes that photon has both a wave and a particle nature, and that these are inseparable.
-Experiments have shown that the above statements are true for any quantum systems including single electron, atom, neutron, and so on.
-In general, since the wave and particle properties of a quantum object coexist simultaneously, one may assume that the two attributes cannot be separated from each other.
-However, recent works on the topic (Chowdhury et al., 2021; Li et al., 2023) has shown, upon exploiting the idea of quantum Cheshire cat (Aharonov et al., 2013), that one may spatially separates of wave and particle attributes of a quantum object.
-The Cheshire cat is a fictitious character in the famous novel “Alice in Wonderland”.
-The Cheshire cat has the ability of disappearing in space while its grin is still visible.
-In our usual life, we can imagine a cat without a grin, but it beyond our common sense to accept the existence of a grin without a cat as Alice said : “Well!
+Measurement problems plays essential roles for our understanding of the nature of the quantum theory and leads to many counterintuitive quantum phenomena.
+One of them is the wave-particle duality of quantum objects which depends on measurement apparatus (Wheeler, 1984).
+In turn, quantum objects display its wave- or particle-like behavior depending on the experimental technique used to measure them (Grangier et al., 1986; Kurtsiefer et al., 1997; DÃŒrr et al., 1998; Marshall et al., 2003; Wang et al., 2019).
+The fundamental issues related to wave-particle duality and Bohr’s complementarity principle (Bohr, 1928; Wootters and Zurek, 1979; Scully et al., 1991) can be exemplified by Young-like double-slit experiments and their variants, e.g.
+Wheeler’s delayed choice experiment (Wheeler, 1978).
+It has been implemented in variety of systems such as photons, atoms, and superconducting circuits (Englert, 1996; Jacques et al., 2007, 2008; Ma et al., 2012; Manning et al., 2015; Tang et al., 2012; Peruzzo et al., 2012; Liu et al., 2017; Ma et al., 2016; Hellmuth et al., 1987; Dong et al., 2020).
+Furthermore, the particle and wave-like behaviors of a single photon (Ionicioiu and Terno, 2011) and multi-entangled photons (Wang et al., 2019) has also been observed simultaneously in one and the same experimental setup (Adesso and Girolami, 2012; Shadbolt et al., 2014).
+Those innovative works provides a new perspective for the in-depth understanding of wave-particle duality.
+In general, since the wave and particle properties of a quantum object coexist simultaneously, one may always suppose to that the two attributes of single entity cannot be separated from each other.
+However, recent research results on the topic (Chowdhury et al., 2021; Li et al., 2023) has shown, upon exploiting the idea of Quantum Cheshire Cat (QCC) (Aharonov et al., 2013), that one may spatially separates of wave and particle attributes of a quantum object.
+The QCC is a fictitious character in the famous novel “Alice in Wonderland”.
+As described in that novel, the Cheshire cat has the ability of disappearing in space while its grin is still visible.
+However, in our usual life, we can imagine a cat without a grin, but it beyond our common sense to accept the existence of a grin without a cat as Alice said : “Well!
 I’ve often seen a cat without a grin, but a grin without a cat!
 It’s the most curious thing I ever saw in my life!”.
-In 2013, Y.
-Aharonov and his collaborators introduced (Aharonov et al., 2013) a protocol to detach a quantum object from its inherent feature, and refer to the result as a quantum Cheshire cat.
-Quantum Cheshire cat may exist in a quantum system described by two-state vector formalism (Aharonov et al., 1964; Aharonov and Vaidman, 1990, 2008).
-Furthermore, its realization depends on performing a post-selected weak measurement (Aharonov et al., 1988) to extract the weak values (Aharonov and Botero, 2005) of the observables associated to the quantum object under investigation.
-In general, the weak value of an observable is a complex number and it can take arbitrary large values (Aharonov and Rohrlich, 2005).
-This feature of the weak value is usually referred to as an amplification effect of post-selection, and can be used to amplify tiny but useful information about physical systems (Hosten and Kwiat, 2008; Dixon et al., 2009; Starling et al., 2009, 2010; Pfeifer and Fischer, 2011; Egan and Stone, 2012; Zhou et al., 2013; Viza et al., 2013; Magaña Loaiza et al., 2014).
-Besides, weak measurement have been also employed in quantum state tomography (Lundeen et al., 2011; Lundeen and Bamber, 2012; de Gosson and de Gosson, 2012; Fischbach and Freyberger, 2012; Matzkin and Pan, 2013; Malik et al., 2013; Salvail et al., 2013; Di Lorenzo, 2013; Malik et al., 2014; Hariri et al., 2019; Pan et al., 2019), Hardy’s paradox (Aharonov et al., 2002; Jiang et al., 2018), three box paradox (Albert et al., 1985; Matzkin and Pan, 2013; Imran et al., 2021; Blasiak and Borsuk, 2022) and so on.
-For more details of weak measurement technique, we refer the reader to recent reviews of the field (Kofman et al., 2012; Svensson, 2013; Sokolovski, 2013; Dressel et al., 2014).
-After the original QCC proposal (Aharonov et al., 2013), the phenomenon has been observed experimentally (Denkmayr et al., 2014) and received large attention theoretically (Guryanova et al., 2012; Matzkin and Pan, 2013; Lorenzo, 2013; Bancal, 2014; Ibnouhsein and Grinbaum, 2014; Corrêa et al., 2015; Duprey et al., 2018; Das and Pati, 2019, 2020; Das and Sen, 2021; Aharonov et al., 2021; Chowdhury et al., 2021; Hance et al., 2023a, b; Zhou et al., 2023; Ghoshal et al., 2023) and experimentally (Atherton et al., 2015; Ashby et al., 2016; Liu et al., 2020; Kim et al., 2021; Wagner et al., 2023; Danner et al., 2023; Li et al., 2023).
-Among the different studies, in (Das and Pati, 2020; Liu et al., 2020) it has been proved that a quantum system may have a physical property erased and may obtain a new one that is did not initially have.
-More recently, a protocol to separate the wave and particle features of a single photon has been proposed (Chowdhury et al., 2021) and implemented (Li et al., 2023).
-Furthermore, in (Zhou et al., 2023), it has been investigated the possibility to spatially separate two entangled photons from their polarizations.
-Since entanglement is a crucial resource for quantum information processing and communication, a question arises on whether the separation of wave and particle attributes for entangled systems is possible.
+The QCC proposal introduced by Y.
+Aharonov and his collaborators (Aharonov et al., 2013) is aim to detach a quantum object from its inherent features.
+QCC cat may exist in a quantum system described by two-state vector formalism (Aharonov et al., 1964; Aharonov and Vaidman, 1990, 2008).
+Furthermore, its realization depends on performing a post-selected weak measurement (Aharonov et al., 1988; Kofman et al., 2012; Svensson, 2013; Sokolovski, 2013; Dressel et al., 2014) to extract the weak values (Aharonov and Botero, 2005) of the observables associated to the quantum object under investigation.
+After the original QCC scheme proposed, it have received large attention and has been investigated theoretically (Guryanova et al., 2012; Matzkin and Pan, 2013; Lorenzo, 2013; Bancal, 2014; Ibnouhsein and Grinbaum, 2014; Corrêa et al., 2015; Duprey et al., 2018; Das and Pati, 2019, 2020; Das and Sen, 2021; Aharonov et al., 2021; Chowdhury et al., 2021; Hance et al., 2023a, b; Zhou et al., 2023; Ghoshal et al., 2023) and experimentally (Atherton et al., 2015; Ashby et al., 2016; Liu et al., 2020; Kim et al., 2021; Wagner et al., 2023; Danner et al., 2023; Li et al., 2023; Denkmayr et al., 2014).
+Among the different studies, a protocol to separate the wave and particle attributes of a single photon has been proposed (Chowdhury et al., 2021) and implemented (Li et al., 2023).
+Furthermore, in (Rab et al., 2017), it has been experimentally observed that two photons can be cast in a wave-particle entangled state provided that suitable initial entangled polarization states are injected into the apparatus.
+Although the separation of particle and wave-like behaviors of a single photon has been observed theoretically and experimentally, the possibility of such observation regarding multi-photon states remains to be explored.
+Since the entanglement is a key resource for quantum information and quantum communication, a question arises on whether the separation of wave- particle attributes of two entangled photons is possible.
 As we will see in the following, our results provide an affirmative answer to this question.
-In this paper, motivated by previous studies (Chowdhury et al., 2021; Li et al., 2023; Zhou et al., 2023), we introduce a feasible theoretical model to spatially separate wave and particle properties of two entangled photons.
-After appropriate pre- and post-selection of the paths of photons, we can find the wave and particle features of the photons in different places.
+In this paper, motivated by previous studies (Rab et al., 2017; Chowdhury et al., 2021; Li et al., 2023), we introduce a feasible theoretical model to spatially separate wave and particle attributes of two entangled photons based on QCC phenomena.
+After appropriate pre- and post-selection of the paths of photons, we can find the wave and particle features of two entangled photons in different places.
 Besides their fundamental interest in the understanding of counterfactual communication, our results may promote the development of quantum metrology.
 The remaining part of this paper is organized as follows.
-In Section II, we illustrate our motivations and briefly introduce the idea of quantum Cheshire cat.
-In SectionIII, we describe in details our proposal and show that by suitable pre- and post-selection one may realize the separation of wave and particle attributes of two entangled photons.
+In Section II,we briefly introduce the idea of quantum Cheshire cat.
+In SectionIII, we describe in details our proposal and show that by suitable pre- and post-selection one can realize the separation of wave and particle attributes of two entangled photons.
 In the same Section, we also describe a possible implementation of our scheme on an optical platform.
-Finally, in Section IV, we close the paper with some concluding remarks and outlooks.
-II The quantum cheshire cat
+Finally, in Section IV, we close the paper with some concluding remarks and discussions.
+II Brief introduction of QCC
 In general, an object and its inherent properties always coexist and cannot be separated at anytime.
-Aharonov and collaborators found that the polarization (say, the grin) property can be separated from a photon (say, the cat) itself in a pre- and post-selected system.
-Here, we briefly summarize the ingredients required to generate such a quantum Cheshire cat.
-Initially, a horizontally polarized photon enters a Mach-Zehnder interferometer (MZI), and the pre-selected state of total system prepared is
-$|\psi_{i}\rangle=\frac{1}{\sqrt{2}}\left(i|L\rangle+|R\rangle\right)|H\rangle,$ (1)
+In original proposal of Aharonov and collaborators (Aharonov et al., 2013), they found that the polarization (say, the grin) property can be separated from a photon (say, the cat) itself in a pre- and post-selected system.
+Here, we briefly summarize the ingredients required to observe such phenomena.
+Initially, a horizontally polarized photon enters a Mach-Zehnder interferometer (MZI) and passes through its right and left arms with equal probability, and the corresponding state reads as
+$|\phi_{i}\rangle=\frac{1}{\sqrt{2}}\left(i|L\rangle+|R\rangle\right)|H\rangle,$ (1)
 where $|L\rangle$ and $|R\rangle$ denote the left and right arms of MZI, i.e.
 the two possible paths of the photon, and $|H\rangle$ denotes the horizontal polarization of the photon.
+We choose the state $|\phi_{i}\rangle$ as our preselected state.
 By adding and adjusting some optical elements, we may post-select the paths and polarization of the photon to obtain the following state vector
-$|\psi_{f}\rangle=\frac{1}{\sqrt{2}}\left(|L\rangle|H\rangle+|R\rangle|V\rangle\right),$ (2)
+$|\phi_{f}\rangle=\frac{1}{\sqrt{2}}\left(|L\rangle|H\rangle+|R\rangle|V\rangle\right),$ (2)
 where $|V\rangle$ represents the vertical polarization of the photon.
@@
 In the two-state-vector formalism, the weak value of an observable $C$ is given by
-$\langle C\rangle_{w}=\frac{\langle\psi_{f}|C|\psi_{i}\rangle}{\;\langle\psi_{f}|\psi_{i}\rangle}.$ (3)
-The location of the photon, i.e, the findings of spatial degree of freedom of the photon, is described by the projectors
+$\langle C\rangle_{w}=\frac{\langle\phi_{f}|C|\phi_{i}\rangle}{\;\langle\phi_{f}|\phi_{i}\rangle}.$ (3)
+This weak value can describe the true value of observable $C$ in time interval corresponded to pre-and postselected states $|\phi_{i}\rangle$ and $|\phi_{f}\rangle$.
+If we want to determine the location of the photon and its polarization in left or right arms of MZI in the region from $|\phi_{i}\rangle$ to $|\phi_{f}\rangle$, we should define the corresponding observables to those quantities.
+Explicitly, we have the operators
 $\Pi_{\mu}=|\mu\rangle\langle\mu|,$ (4)
-where $\mu\in\left(L,R\right)$, and $\Pi_{L}\left(\Pi_{R}\right)$ denote left (right) arm of MZI.
-On the other hand, the circular polarization of the photon in the left or right arms of MZI can be characterized by the operator,
 $\sigma_{z}^{\mu}=\sigma_{z}\otimes\Pi_{\mu}.$ (5)
-where $\sigma_{z}=|\uparrow_{y}\rangle\langle\uparrow_{y}|-|\downarrow_{y}\rangle\langle\downarrow_{y}|$ with $|\uparrow_{y}\rangle=\frac{1}{\sqrt{2}}\left(|H\rangle+i|V\rangle\right)$ and $|\downarrow_{y}\rangle=\frac{1}{\sqrt{2}}\left(|H\rangle-i|V\rangle\right)$, i.e.
-the eigenstate of the Pauli matrix $\sigma_{y}=\left(\begin{array}[]{cc}0&\ -i\\ i&0\end{array}\right)$ with corresponding eigenvalues $1$ and $-1$).
-By taking into account the above pre- and post-selected states $|\psi_{i}\rangle$ and $|\psi_{f}\rangle$, the weak value of $\Pi_{\mu}$ and $\sigma_{z}^{\mu}$ are given by
+Here, $\Pi_{L}\left(\Pi_{R}\right)$ with $\mu\in\left(L,R\right)$ is the projection operator of left (right) arm of MZI and $\sigma_{z}=|\uparrow_{y}\rangle\langle\uparrow_{y}|-|\downarrow_{y}\rangle\langle\downarrow_{y}|$ is circular polarization observable of the photon with $|\uparrow_{y}\rangle=\frac{1}{\sqrt{2}}\left(|H\rangle+i|V\rangle\right)$ and $|\downarrow_{y}\rangle=\frac{1}{\sqrt{2}}\left(|H\rangle-i|V\rangle\right)$.
+The weak values of $\Pi_{\mu}$ and $\sigma_{z}^{\mu}$ can characterize the definite values of exact location of spatial degree of freedom and circular polarization of the photon inside the MZI, respectively.
+By taking into account the above pre- and post-selected states $|\phi_{i}\rangle$ and $|\phi_{f}\rangle$, the weak value of $\Pi_{\mu}$ and $\sigma_{z}^{\mu}$ are given by
 $\langle\Pi_{\mu}\rangle_{w}=\delta_{\mu L},$ (6)
@@
 $\langle\sigma_{z}^{\mu}\rangle_{w}=\delta_{\mu R},$ (7)
-respectively. Here, $\delta_{ij}$ denote the Kronecker $\delta$ function which defined as
+respectively. Here, $\delta_{ij}$ denote the Kronecker $\delta$-function which defined as
 $\delta_{ij}=\begin{cases}1,&i=j\\ 0.&i\neq j\end{cases}$ (8)
+We can see that $\langle\Pi_{L}\rangle_{w}=\langle\sigma_{z}^{R}\rangle=1$ and other weak values all equal to zero.
 The weak values in Eq.(6) and Eq.
-(7) indicates that the spatial degree freedom of the photon is located in the left arm, while its circular polarization is detected with unit probability in the right arm of the MZI.
-This mean that in pre- and post-selected quantum systems one can realize the separation of the object from its inherent properties, as the Cheshire cat, whose smile may be discovered at a different place of the cat itself.
+(7) indicated that the spatial degree freedom of the photon is located in the left arm (photon itself go through the left arm of MZI), while its circular polarization is detected with unit probability in the right arm of the MZI.
+This means that in properly pre- and post-selected quantum systems one can realize the separation of the object itself from its inherent properties, as the Cheshire cat.
 III Separation of wave-particle properties of two entangled photons
-In this section, we describe our proposal for the separation of wave-particle properties of two-entangled photons.
+In this section, we describe our proposal for the separation of wave-particle properties of two entangled photons.
 The schematic diagram of our proposed setup is shown in Fig.
 1.
-It is based on the so-called WP toolbox introduced in (Rab et al., 2017) to convert the coherence superposition of the polarization states to the coherence superposition of the wave and particle entities.
-Figure 1: Schematic diagram of the scheme to spatially separate the wave and particle properties of a two-photon entangled state.
-In (Rab et al., 2017) single-photon states and two-photon entangled states have been analyzed.
-The generation processes of the wave-particle entangled state of two separated photons can be treated as parallel doubling of the single-photon scheme.
-In particular, upon using a half-wave plate (HWP, not shown in the figure) on the path of both the photons produced by spontaneous parametric down-conversion (SPDC), the two-photon state becomes
+The one of most important points of our scheme is the preparation of so-called wave-particle (WP) toolbox introduced in Ref.
+(Rab et al., 2017) to convert superpositions of polarization states to the superpositions of path and polarization states.
+Next we describe our scheme step by step along with our schematic diagram depicted in Fig.
+1.
+Figure 1: Illustration of separating the wave and particle attributes of two entangled photons.
+We assume that in type-I spontaneous parametric down-conversion (SPDC) process two photons with same polarization have generated, i.e.
+$|V\rangle_{1}|V\rangle_{2}$ or $|H\rangle_{1}|H\rangle_{2}$.
+Upon using a half-wave plate (HWP, not shown in the figure) rotated at an angle $\alpha/2$ on the path of both the photons, the two-photon state becomes
 $|\psi\rangle_{in}=\cos\alpha|V\rangle|V\rangle+\sin\alpha|H\rangle|H\rangle,$ (9)
-where $\alpha$ is an adjustable parameter, tunable by the HWP. After this stage, each photon enters one of two identical WP toolboxes, and the final output state may be written as (Rab et al., 2017)
-$|\Phi\rangle_{out}=\cos\alpha|W\rangle|W^{\prime}\rangle+\sin\alpha|P\rangle|P^{\prime}\rangle.$ (10)
-Here,
-$|W\rangle$ $\equiv|Wave\rangle=e^{i\frac{\phi_{1}}{2}}\left(\cos\frac{\phi_{1}}{2}|1\rangle-i\sin\frac{\phi_{1}}{2}|3\rangle\right),$ (11)
-$|W^{\prime}\rangle$ $\equiv\!\!|Wave^{\prime}\rangle=e^{i\frac{\phi_{1}}{2}}\left(\cos\frac{\phi_{1}}{2}|1^{\prime}\rangle\!-i\sin\frac{\phi_{1}}{2}|3^{\prime}\rangle\right),$ (12)
-$|P\rangle$ $\equiv|Particle\rangle=\frac{1}{\sqrt{2}}\left(|2\rangle+e^{i\phi_{1}}|4\rangle\right),$ (13)
-$|P^{\prime}\rangle$ $\equiv|Particle\rangle=\frac{1}{\sqrt{2}}\left(|2^{\prime}\rangle+e^{i\phi_{1}}|4^{\prime}\rangle\right).$ (14)
-In the above expressions, $|n\rangle$ and $|n^{\prime}\rangle$ with $n,n^{\prime}\in\{1,2,3,4\}$ being the $n$-th output modes of from the two WP toolboxes (Rab et al., 2017), and $\phi_{1}$, $\phi_{1}^{\prime}$, $\phi_{2}$ and $\phi_{2}^{\prime}$ being controllable phase shifts in the two WP toolboxes.
-For more details about this generation of the states in Eqs.
-(11-14) we refer to the original paper (Rab et al., 2017).
-Notice that these states represent the capabilities ($|W\rangle$and $|W^{\prime}\rangle$) and incapabilities ($|P\rangle$ and $|P^{\prime}\rangle$) of the two photons to produce interference.
+where $\alpha$ is an adjustable parameter. After this stage, each photon enters one of two identical WP toolboxes, and the final output state can be written as (Rab et al., 2017)
+$|\Phi\rangle_{out}$ $=\cos\alpha|W\rangle_{1}|W^{\prime}\rangle_{2}+\sin\alpha|P\rangle_{1}|P^{\prime}\rangle_{2}$
+$=\cos\alpha|W\rangle|W^{\prime}\rangle+\sin\alpha|P\rangle|P^{\prime}\rangle,$ (10)
+where
+$|W\rangle$ $\equiv|Wave\rangle=e^{i\frac{\phi_{1}}{2}}\left(\cos\frac{\phi_{1}}{2}|1\rangle-i\sin\frac{\phi_{1}}{2}|3\rangle\right),$ (11a)
+$|W^{\prime}\rangle$ $\equiv\!\!|Wave^{\prime}\rangle=e^{i\frac{\phi_{1}}{2}}\left(\cos\frac{\phi_{1}}{2}|1^{\prime}\rangle\!-i\sin\frac{\phi_{1}}{2}|3^{\prime}\rangle\right),$ (11b)
+$|P\rangle$ $\equiv|Particle\rangle=\frac{1}{\sqrt{2}}\left(|2\rangle+e^{i\phi_{1}}|4\rangle\right),$ (11c)
+$|P^{\prime}\rangle$ $\equiv|Particle\rangle=\frac{1}{\sqrt{2}}\left(|2^{\prime}\rangle+e^{i\phi_{1}}|4^{\prime}\rangle\right),$ (11d)
+with $|n\rangle$ and $|n^{\prime}\rangle$ ( $n,n^{\prime}\in\{1,2,3,4\}$ ) being the $n(n^{\prime})$-th output modes of from the two WP toolboxes (Rab et al., 2017), and $\phi_{1}$, $\phi_{1}^{\prime}$, $\phi_{2}$ and $\phi_{2}^{\prime}$ being controllable phase shifts in the two WP toolboxes.
+Notice that these states can characterize the capabilities ($|W\rangle$and $|W^{\prime}\rangle$) and incapabilities ($|P\rangle$ and $|P^{\prime}\rangle$) of the two photons to produce interference.
+In (Rab et al., 2017), the generation processes of superpositions of wave and particle states for single photon and two entangled photons both have been analyzed.
+For more details about the generation of the state, i.g, Eqs.
+(11a-11d) in the lab we refer to the original paper (Rab et al., 2017).
 Finally, as shown in Fig.
@@
 This is our pre-selected state.
-Here, as quantum Cheshire cat’s notation, we use the composite system of path and attributes of two photons as |..$|.\ .\rangle_{path}|.\ .\rangle_{attributes}$.
-$|L_{i}\rangle$ and $|R_{i}\rangle$ denoting the left and right paths , $i\in\{1,2\}$ labels the photons.
+Here, as the notation of QCC, we use the composite system of path and attributes of two photons as |..$|.\ .\rangle_{paths}|.\ .\rangle_{attributes}$.
+$|L_{i}\rangle$ and $|R_{i}\rangle$ denoting the left and right paths, and $i=1,2$ labels the first and second photons.
 $|R_{1}L_{2}\rangle|WW^{\prime}\rangle$ represents $1$st (2nd) photon’s wave property $|W\rangle$ ($|W^{\prime}\rangle$) at $|R_{1}\rangle$ ($|L_{2}\rangle$), and $|L_{1}R_{2}\rangle|PP{}^{\prime}\rangle$ represents $1$st (2nd) photon’s particle property $|P\rangle$ ($|P^{\prime}\rangle$) at $|L_{1}\rangle$ ($|R_{2}\rangle$).
-$|W\rangle$, $|W^{\prime}\rangle$ and $|P\rangle$, $|P^{\prime}\rangle$ denote, respectively, the wave and particle states as (Ionicioiu and Terno, 2011; Rab et al., 2017)
 In order to realize the separation of wave and particle attributes of two entangled photons, the post-selected state should be of the form
@@
 This means that in our scheme the post-selected state is a special case of the pre-selected class of states with $\alpha=\frac{\pi}{4}$.
-The determination of the post-selected state $|\psi_{f}\rangle$ is crucial for the implementation of our scheme.
+The determination of post-selected state $|\psi_{f}\rangle$ is crucial for the implementation of our scheme.
 As shown in Fig.
-1, its verification can accomplished by an optical setup made of five beam splitters $BS_{i}$ ($i=1,2,3,4,5$), two $\sigma^{1224}$operators, the $X_{3}$ and $X_{4}$ operators, three mirrors $M_{i}$ ($i=1,2,3$), and six detectors $D_{i}$ ($i=1,2,3,4,5,6$).
-The post-selection occurs if and only if the detector $D_{5}$ clicks.
-On the contrary, if our post-seleceted state is $|\psi_{f}\rangle$, then, in the optical arrangement of Fig.
-1, we can guarantee only $D_{5}$ give the answer ,“yes”.
+1, its verification can accomplished by an optical setup made of five beam splitters $BS_{i}$ ($i=1,2,3,4,5$), two $\sigma^{1234}$ operators, two $X_{i}$ operators, three mirrors $M_{i}$ ($i=1,2,3$), and six detectors $D_{i}$ ($i=1,2,3,4,5,6$).
+The successful post-selection occurs iff the detector $D_{5}$ click.
+On the contrary, if our post-seleceted state to be in $|\psi_{f}\rangle$, then, in the optical arrangement of Fig.
+1, we can guarantee only the detector $D_{5}$ give the answer ,“yes”.
 Let us prove this statement.
-At first, the photons in $L_{1}$ and $R_{2}$ arms go through the $BS$ followed by $\sigma^{1234}$ operator.
-Here, effect of $\sigma^{1234}$ operator is to switch the photon paths in the four-dimensional basis $\{|1\rangle,|2\rangle,|3\rangle,|4\rangle\}$.
+At first, the photons in $L_{1}$ and $R_{2}$ arms simultaneously go through the beam splitters $BS_{1}$ and $BS_{2}$ followed by $\sigma^{1234}$ operators.
+Here, the effect of $\sigma^{1234}$ operator is to switch the photon paths in the four-dimensional basis $\{|1\rangle,|2\rangle,|3\rangle,|4\rangle\}$.
 Its matrix form can be written as (Rab et al., 2017)
@@
 $|1\rangle\stackrel{{\scriptstyle[}}{{\sigma}}^{1234}]{\sigma^{1234}}{\rightleftharpoons}|2\rangle,\ |3\rangle\stackrel{{\scriptstyle[}}{{\sigma}}^{1234}]{\sigma^{1234}}{\rightleftharpoons}|4\rangle.$ (16)
-The beam splitter $BS_{1}$ is chosen as to perform the transformations $|2\rangle\rightarrow\frac{1}{\sqrt{2}}(|2\rangle+|4\rangle)$, and $|4\rangle\rightarrow\frac{1}{\sqrt{2}}(|2\rangle-|4\rangle)$.
-In this way, the particle state can converted to the wave state after the action of $BS_{1}$ and $\sigma^{1234}$, i.e,
+The beam splitters $BS_{1}$ and $BS_{2}$ are adjusted to perform the transformations as
+$|2\rangle$ $\overset{BS_{1}}{\longrightarrow}\frac{1}{\sqrt{2}}(|2\rangle+|4\rangle)$ (17a)
+$|4\rangle$ $\overset{BS_{1}}{\longrightarrow}\frac{1}{\sqrt{2}}(|2\rangle-|4\rangle)$ (17b)
+and
+$|2^{\prime}\rangle$ $\overset{BS_{1}}{\longrightarrow}\frac{1}{\sqrt{2}}(|2^{\prime}\rangle+|4^{\prime}\rangle)$ (17c)
+$|4^{\prime}\rangle$ $\overset{BS_{1}}{\longrightarrow}\frac{1}{\sqrt{2}}(|2^{\prime}\rangle-|4^{\prime}\rangle)$ (17d)
+In this way, the particle state can converted to the wave state after the action of $BS_{1}$ ($BS_{2}$) and $\sigma^{1234}$, i.e,
 $|P\rangle\overset{BS_{1}}{\longrightarrow}$ $\frac{1}{\sqrt{2}}\left[\frac{1}{\sqrt{2}}(|2\rangle+|4\rangle)+e^{i\phi_{1}}\frac{1}{\sqrt{2}}(|2\rangle-|4\rangle)\right]$
@@
 $e^{i\frac{\phi_{1}}{2}}\left(\cos\frac{\phi_{1}}{2}|1\rangle-i\sin\frac{\phi_{1}}{2}|3\rangle\right)=|W\rangle.$ (18)
-The same happens for primed paths with $|n^{\prime}\rangle\in\left(|1^{\prime}\rangle,|2^{\prime}\rangle,|3^{\prime}\rangle,|4^{\prime}\rangle\right)$.
-Thus, after the photons in $L_{1}$ and $R_{2}$ arms simultaneously go through beam splitters $BS_{1}$ and $BS_{2}$ followed by the $\sigma^{1234}$ operators, our post-selected state is given by
+The same happens for primed paths with $|n^{\prime}\rangle\in\left(|1^{\prime}\rangle,|2^{\prime}\rangle,|3^{\prime}\rangle,|4^{\prime}\rangle\right)$, as
+$|P^{\prime}\rangle\overset{BS_{1}}{\longrightarrow}e^{i\frac{\phi_{1}^{\prime}}{2}}\left(\cos\frac{\phi_{1}^{\prime}}{2}|2^{\prime}\rangle-i\sin\frac{\phi_{1}^{\prime}}{2}|4^{\prime}\rangle\right)\overset{\sigma^{1234}}{\longrightarrow}|W^{\prime}\rangle$ (19)
+Thus, after the photons in $L_{1}$ and $R_{2}$ arms simultaneously go through beam splitters $BS_{1}$ and $BS_{2}$ followed by the $\sigma^{1234}$ operators, our post-selected state $|\psi_{f}\rangle$ changed to
 $|\psi_{f^{\prime}}\rangle=\frac{1}{\sqrt{2}}\left(|R_{1}L_{2}\rangle+|L_{1}R_{2}\rangle\right)|WW^{\prime}\rangle.$ (20)
-Next, by adjusting the beam-splitters $BS_{3}$ and $BS_{4}$, respectively, we make the states $|R_{1}L_{2}\rangle$ and $|L_{1}R_{2}\rangle$ go through $X_{3}$ and $X_{4}$, such that the detectors $D_{1}$ and $D_{2}$ will not click.
+Next, by adjusting the beam-splitters $BS_{3}$ and $BS_{4}$, respectively, such that states $|R_{1}\rangle$ and $|L_{1}\rangle$ ($|L_{2}\rangle$ and $|R_{2}\rangle$) totally transmitted (reflected) to guarantee $|R_{1}L_{2}\rangle$ and $|L_{1}R_{2}\rangle$ go through $X_{3}$ and $X_{4}$.
+Under this action the detectors $D_{1}$ and $D_{2}$ both will not click.
+In this step, the modified form of $|\psi_{f^{\prime}}\rangle$ can be written as
+$|\psi_{f^{\prime}}\rangle\overset{BS_{3},BS_{4}}{\longrightarrow}|\psi_{f^{\prime\prime}}\rangle=\frac{1}{\sqrt{2}}\left(|R\rangle+|L\rangle\right)|\mathcal{\boldsymbol{W}}\rangle.$ (21)
+Here, we denote $|R_{1}L_{2}\rangle\overset{BS_{3}}{\longrightarrow}|R\rangle$, $|L_{1}R_{2}\rangle\overset{BS_{4}}{\longrightarrow}|L\rangle$ and $|\mathcal{\boldsymbol{W}}\rangle=|WW^{\prime}\rangle$.
 The $X_{3}=X_{4}=|W\rangle\langle W|$ operators transmit only the wave state, and this action guarantees that detectors $D_{3}$ and $D_{4}$ will not click.
-Finally, by adjusting the beam splitter $BS_{5}$, the state $|R_{1}L_{2}\rangle+|L_{1}R_{2}\rangle$ reaches the detector $D_{5}$.
-Overall, we conclude that if and only if the detector $D_{5}$ clicks we have made a successful post-selection.
-Otherwise, if any other detector clicks post-selection fails, and we fail to separate the wave and particle attributes of the two entangled photons.
+The beam splitter $BS_{5}$ is chosen as $|R\rangle\overset{BS_{5}}{\longrightarrow}\frac{1}{\sqrt{2}}\left(|R\rangle+|L\rangle\right)$, $|L\rangle\overset{BS_{5}}{\longrightarrow}\frac{1}{\sqrt{2}}\left(|R\rangle-|L\rangle\right)$, such that
+$|\psi_{f^{\prime\prime}}\rangle\overset{BS_{5}}{\longrightarrow}|\Psi_{f}\rangle=|R\rangle|\mathcal{\boldsymbol{W}}\rangle.$ (22)
+Hence, we can confirm that detector $D_{5}$ is click with certainty.
+Contrarily, if and only if the detector $D_{5}$ click with $100\%$ probability, then our schematics also can prove that the postslected state should be chosen to $|\psi_{f}\rangle$.
+After choosing the pre- and post-selected state of our scheme, we have to extract the weak values to achieve our goal.
 The operators that measure whether the wave and particle attributes of two entangled photons are present in the left and right arms is defined as
 $\Pi_{\nu}^{\mu_{i}}=|\mu_{i}\rangle\langle\mu_{i}|\otimes|\nu\rangle\langle\nu|,$ (23)
-with $\nu\in\{W,W^{\prime},P,P^{\prime}\}$, $\mu_{i}\in\{L_{i},R_{i}\}$ and $i\in\{1,2\}$.
-After preparing the pre-selected state, we choose to post-select by $|\psi_{f}\rangle$ and perform weak measurement for observables $\Pi_{\nu}^{\mu_{i}}$.
-By substituting the above-defined pre- and post-selected states into Eq.
-(3), we find the corresponding weak values of the observables $\Pi_{\nu}^{\mu_{i}}$, i.e.,
-$\langle\Pi_{W}^{\mu_{i}}\rangle$ $=\frac{\cos\alpha}{\cos\alpha+\sin\alpha}\delta_{1i}\delta_{R_{i}\mu_{i}},$ (23a)
-$\langle\Pi_{P}^{\mu_{i}}\rangle$ $=\frac{\sin\alpha}{\cos\alpha+\sin\alpha}\delta_{1i}\delta_{L_{i}\mu_{i}},$ (23b)
-$\langle\Pi_{W^{\prime}}^{\mu_{i}}\rangle$ $=\frac{\cos\alpha}{\cos\alpha+\sin\alpha}\delta_{2i}\delta_{L_{i}\mu_{i}},$ (23c)
-$\langle\Pi_{P^{\prime}}^{\mu_{i}}\rangle$ $=\frac{\sin\alpha}{\cos\alpha+\sin\alpha}\delta_{2i}\delta_{R_{i}\mu_{i}}.$ (23d)
-If $\alpha\neq 0,\frac{\pi}{2}$, these weak values indicated that the wave and particle attributes of two entangled photons are perfectly separated, and in each one of the four arms we only have one attribute (wave or particle ) of the two photons.
-In the particular case, $\alpha=\frac{\pi}{4},$ the weak values for the wave and particle attributes may be written as $\langle\Pi_{P}^{L_{1}}\rangle=\langle\Pi_{W}^{R_{1}}\rangle=\frac{1}{2}$, $\langle\Pi_{P^{\prime}}^{R_{2}}\rangle=\langle\Pi_{W^{\prime}}^{L_{2}}\rangle=\frac{1}{2}$.
-This means that for $\alpha=\frac{\pi}{4}$ case, half of the particle attributes of the two entangled photons occurrs in $L_{1}$ and $R_{2}$ arms and half of the wave attributes are present in $R_{1}$ and $L_{2}$ arms of the interferometer, respectively.
+with $\nu\in\{W,W^{\prime},P,P^{\prime}\}$, $\mu_{i}\in\{L_{i},R_{i}\}$ with $i=1,2$.
+Above every operators has its meaning, such as $\Pi_{W}^{L_{1}}=|L_{1}\rangle\langle L_{1}|\otimes|W\rangle\langle W|$ can describe the findings of the wave attribute of first photon on the left arm and $\Pi_{P^{\prime}}^{R_{2}}=|R_{2}\rangle\langle R_{2}|\otimes|P^{\prime}\rangle\langle P^{\prime}|$ can describe the findings of the particle attribute of second photon on the right arm, etc.
+By substituting the above-defined pre- and post-selected states $|\psi_{f}\rangle$ and $|\psi_{f}\rangle$ into Eq.
+(3), we obtain the corresponding weak values of the observables $\Pi_{\nu}^{\mu_{i}}$, i.e.,
+$\langle\Pi_{W}^{\mu_{i}}\rangle_{w}$ $=\frac{\cos\alpha}{\cos\alpha+\sin\alpha}\delta_{1i}\delta_{R_{i}\mu_{i}},$ (24a)
+$\langle\Pi_{P}^{\mu_{i}}\rangle_{w}$ $=\frac{\sin\alpha}{\cos\alpha+\sin\alpha}\delta_{1i}\delta_{L_{i}\mu_{i}},$ (24b)
+$\langle\Pi_{W^{\prime}}^{\mu_{i}}\rangle_{w}$ $=\frac{\cos\alpha}{\cos\alpha+\sin\alpha}\delta_{2i}\delta_{L_{i}\mu_{i}},$ (24c)
+$\langle\Pi_{P^{\prime}}^{\mu_{i}}\rangle_{w}$ $=\frac{\sin\alpha}{\cos\alpha+\sin\alpha}\delta_{2i}\delta_{R_{i}\mu_{i}}.$ (24d)
+This is the main result of our study, and they can characterize the exact values of operators $\Pi_{\nu}^{\mu_{i}}$ in the region from $|\psi_{i}\rangle$ to $|\psi_{f}\rangle$.
+If $\alpha\neq 0,\frac{\pi}{2}$, these weak values indicated that the wave and particle attributes of two entangled photons are spatially separated, and in each one of the four arms we only have one attribute (wave or particle ) of the two photons.
+In the particular case, $\alpha=\frac{\pi}{4},$ the weak values for the wave and particle attributes may be written as $\langle\Pi_{P}^{L_{1}}\rangle_{w}=\langle\Pi_{W}^{R_{1}}\rangle_{w}=\frac{1}{2}$, $\langle\Pi_{P^{\prime}}^{R_{2}}\rangle_{w}=\langle\Pi_{W^{\prime}}^{L_{2}}\rangle_{w}=\frac{1}{2}$.
+This means that for $\alpha=\frac{\pi}{4}$ case, half of the particle attributes of the two entangled photons occurred in $L_{1}$ and $R_{2}$ arms and half of the wave attributes are present in $R_{1}$ and $L_{2}$ arms of the interferometer, respectively.
 Furthermore, it is clear that these weak values cannot take arbitrary large values, and always obey the Bohr’s complementarity principle, i.e.,
-$\langle\Pi_{P}^{L_{1}}\rangle+\langle\Pi_{W}^{R_{1}}\rangle$ $=1,$ (24)
-$\langle\Pi_{P^{\prime}}^{R_{2}}\rangle+\langle\Pi_{W^{\prime}}^{L_{2}}\rangle$ $=1.$ (25)
-IV conclusion
+$\langle\Pi_{P}^{L_{1}}\rangle_{w}+\langle\Pi_{W}^{R_{1}}\rangle_{w}$ $=1,$ (25)
+$\langle\Pi_{P^{\prime}}^{R_{2}}\rangle_{w}+\langle\Pi_{W^{\prime}}^{L_{2}}\rangle_{w}$ $=1.$ (26)
+In this proposal, the correctly choosing of post-selected state is very crucial and we only focus on the cases of $D_{5}$ click with certainty.
+Otherwise, if any other detector clicks, the post-selection fails, and we fail to separate the wave and particle attributes of two entangled photons.
+IV CONCLUSION AND DISCUSSION
 In this paper, we have investigated the possibility of spatially separating the wave and particle features of two entangled photons.
-Our work is based on measuring the weak values of two-level quantum systems by selecting appropriate pre- and post-selection states in order to generate polarization-path like quantum Cheshire cats.
-We found that properly pre- and post-selecting the two-photon state, the wave and particle attributes of two entangled photons can be spatially separated, while keeping the validity of Bohr’s complementarity principle.
-We also proposed a feasible setup for the implementation of our thought experiment on a quantum optical platform.
-The quantum entanglement is a cornerstone of quantum informatics and it has an irreplaceable role in quantum precision measurements.
+Our work is based on measuring the weak values of system by selecting appropriate pre- and post-selection states in order to generate polarization-path quantum Cheshire cat states of two entangled photons.
+We found that by properly choosing the pre- and post-select states of system, the wave and particle attributes of two entangled photons can be spatially separated, while keeping the validity of Bohr’s complementarity principle.
+We also proposed a feasible setup for the implementation of our thought experiment with quantum optical platform.
+As indicated in our result, we can control wave-particle attributes of two entangled photons by adjusting the optical axis angle $\alpha$ of HWP.
+If $\alpha=0$ or $\alpha=\frac{\pi}{2}$, then we only have pure wave attributes or pure particle attributes of two photons and separate them spatially.
+Thus, it is very curious issue to do double- slit experiment by using pure particle attributes of entangled photons (Hong and Noh, 1998).
+Furthermore, in the emerging field of counterfactual communication (Salih et al., 2013; Li et al., 2015), information may be exchanged even without the transmission of quantum systems.
+In this framework, entanglement also plays a vital role, but the basic principles of such communication is still unclear.
+Thus, it may interesting to use pure wave attributes of entangled photons for understanding of the fundamental mechanism behind counterfactual communication.
+At the same time, our results may also provide an alternative method to enhance the precision of quantum metrological protocols.
 In particular, in quantum metrology one would generally require that the measurement of a physical quantity on an entangled state does non disturb to other variable.
-Our results may foster the exploration of schemes to separate entangled observables and enhance precision of quantum metrological protocols.
-At the same time, in the emerging field of counterfactual communication (Salih et al., 2013; Li et al., 2015), information may be exchanged even without the transmission of quantum state.
-In this framework, entanglement also plays a vital role, but the basic principles of communication is still unclear.
-Also in this case, our proposal may promote the understanding of the fundamental mechanism behind counterfactual communication.
+Thus, the separation of entangled observables by QCC and implement the related precision measurement may dramatically increase its efficiency.
+Another interesting point of our scheme is that it can scale up to more multi-entangled photons cases.
+Since the parallel assembly of $N$ single-photon WP toolboxes allows us to generate $N$-photon wave-particle entangled states (Rab et al., 2017), we may realize the spatial separation of wave and particle attributes of $N$-photon wave-particle entangled states by expand our schematics.
 Finally, we anticipate that our theoretical proposal is feasible with current technology and may be implemented in quantum optical labs in the near future.
 Acknowledgements.
-This work was supported by the National Natural Science Foundation of China (No. 12365005).
+This work was supported by the National Natural Science Foundation of China (No. 12365005)